# Fact-Checking the Output of Large Language Models via Token-Level Uncertainty Quantification

### v1

@@ Appendix F FactScore Annotation System
Here we give examples of the operation of the FactScore automatic markup system, see the Table 9. We hypothesize that the causes of FactScore errors are related to model hallucination (true information is present in the knowledge source, but the model produces incorrect information), lack of context (an excerpt from a Wikipedia article cannot capture all the details), and difficulties with information interpretation (the desired information is present in the knowledge base, but is formulated in different words or in several sentences). All these cases leave room for further improvement of the pipeline. We have also given the results of the proposed CCP method on selected examples. It can be seen that the results of our algorithm are similar to the annotation from FactScore, while CCP does not use external information in calculating the scores.
@@ Appendix G Examples of Results Obtained with CCP and the Maximal Probability Baseline
Figure 10: Claim-Conditioned Probability vs Maximum Probability baseline.
@@
Figure 11: Claim-Conditioned Probability vs Maximum Probability baseline.
@@
Figure 12: Claim-Conditioned Probability vs Maximum Probability baseline.
Appendix H Resources and Expenses
For a single run of data generation and UQ methods evaluation, we spent 12 days of Nvidia A100 GPU compute time. OpenAI API was used mainly for splitting and matching atomic claims, the total cost for all English models (Mistral 7b, Vicuna 13b, Jais 13b, GPT-3.5-turbo) was 40\$. The cost of generating English biographies with GPT-3.5-turbo was 13\$.

### v2

@@ C.3 FactScore Annotation
The statistics of English data annotated using FactScore is presented in Table 7. Here we give examples of the operation of the FactScore automatic markup system, see the Table 8. We hypothesize that the causes of FactScore errors are related to model hallucination (true information is present in the knowledge source, but the model produces incorrect information), lack of context (an excerpt from a Wikipedia article cannot capture all the details), and difficulties with information interpretation (the desired information is present in the knowledge base, but is formulated in different words or in several sentences). All these cases leave room for further improvement of the pipeline. We have also given the results of the proposed CCP method on selected examples. It can be seen that the results of our algorithm are similar to the annotation from FactScore, while CCP does not use external information in calculating the scores.
@@
Appendix D Ablation Studies
In this section, we perform the ablation study of various CCP components.
D.1 Aggregation
Method ROC-AUC PR-AUC
$CCP_{\textit{prod}}$ 0.66 $\pm$ 0.03 0.22 $\pm$ 0.05
$CCP_{\textit{len}}$ 0.65 $\pm$ 0.03 0.21 $\pm$ 0.03
$CCP_{\textit{min}}$ 0.64 $\pm$ 0.03 0.13 $\pm$ 0.04
$CCP_{\textit{mean}}$ 0.65 $\pm$ 0.03 0.22 $\pm$ 0.04
Table 9: ROC-AUC and PR-AUC on Vicuna 13b generation, for different normalizations.
$CCP_{\textit{prod}}(F)$ $=1-\prod_{j=1}^{k}CCP_{\textit{word}}(x_{j}),$
$CCP_{\textit{len}}(F)$ $=1-\exp\biggl(\frac{1}{k}\sum_{j=1}^{k}\log CCP_{\textit{word}}(x_{j})\biggr),$
$CCP_{\textit{min}}(F)$ $=1-\min_{j=1,\dots,k}CCP_{\textit{word}}(x_{j}),$
$CCP_{\textit{mean}}(F)$ $=1-\frac{1}{k}\sum_{j=1}^{k}CCP_{\textit{word}}(x_{j}).$
$CCP_{\textit{prod}}$ shows the best results and is chosen for the final $CCP$ formula.
D.2 NLI Models
NLI model Parameters ROC-AUC PR-AUC
microsoft/deberta-large-mnli 350M 0.66 $\pm$ 0.06 0.24 $\pm$ 0.04
microsoft/deberta-base-mnli 86M 0.65 $\pm$ 0.06 0.23 $\pm$ 0.05
MoritzLaurer/mDeBERTa-v3-base 86M 0.66 $\pm$ 0.06 0.21 $\pm$ 0.03
cross-encoder/nli-deberta-v3-large 350M 0.66 $\pm$ 0.06 0.22 $\pm$ 0.04
cross-encoder/nli-deberta-v3-base 86M 0.65 $\pm$ 0.06 0.23 $\pm$ 0.05
cross-encoder/nli-deberta-v3-small 44M 0.65 $\pm$ 0.06 0.22 $\pm$ 0.05
cross-encoder/nli-deberta-v3-xsmall 22M 0.66 $\pm$ 0.07 0.24 $\pm$ 0.06
Table 10: The ROC-AUC and PR-AUC metrics of CCP on biographies generation dataset with Vicuna 13b model, when using different NLI models with specified identifiers in Huggingface library.
Table 10 presents the result of CCP on Vicuna 13b when using different NLI models. We selected fine-tuned NLI models from HuggingFace, encompassing a range of sizes from Microsoft’s DeBERTa model He et al. (2021), a multilingual variant Laurer et al. (2022), and a model from SentenceTransformers Reimers and Gurevych (2019).
Our findings demonstrate that the CCP performance exhibits minimal dependence on the specific NLI model chosen. Notably, even CCP utilizing a relatively small CrossEncoder model with only 22M parameters achieves the highest performance.
D.3 NLI Context
Method ROC-AUC PR-AUC
$CCP_{\textit{no context}}$ 0.64 $\pm$ 0.03 0.24 $\pm$ 0.04
$CCP_{\textit{sent pref}}$ 0.59 $\pm$ 0.03 0.18 $\pm$ 0.04
$CCP_{\textit{claim pref}}$ 0.66 $\pm$ 0.03 0.24 $\pm$ 0.05
Table 11: ROC-AUC and PR-AUC on Vicuna 13b generation, for different contexts to input words with to an NLI model.
  1. 1.
  2. 2.
  3. 3.
The results for sentence prefix provides too wide context to the NLI model, which results in the NLI model focusing on the contexts instead of the greedy words and its alternatives, thus outputting lots of entailment classes. On the other hand, the method without context does not provide enough context to calculate NLI class with more quality. $CCP_{\textit{claim pref}}$ shows better results and is the one chosen for the final $CCP$ formula.
D.4 Functional Words Handling
Method ROC-AUC PR-AUC
$CCP_{\textit{confident}}$ 0.66 $\pm$ 0.03 0.24 $\pm$ 0.05
$CCP_{\textit{ignore}}$ 0.63 $\pm$ 0.03 0.23 $\pm$ 0.06
$MP_{\textit{confident}}$ 0.59 $\pm$ 0.05 0.18 $\pm$ 0.04
$MP_{\textit{ignore}}$ 0.60 $\pm$ 0.05 0.19 $\pm$ 0.05
Table 12: ROC-AUC and PR-AUC on Vicuna 13b generation, for methods of handling functional words.
  1. 1.
  2. 2.
D.5 Number of Alternatives
Figure 7: ROC-AUC between FactScore classes and the claim-level CCP method as a function of $n$ – the number of token options in the probability distribution of the LLM to consider.
Appendix E Additional Experimental Results
Here we show additional results related to the performance of our pipeline and CCP method. In the Figure 8 we show the dependence of the percentage of supported claims in the generated LLM response (in this case for Vicuna 13b) as a function of its length. As we can see, the LLM produces wrong claims when the generation length increases, which may be due to the generation of additional facts.
In Figure 9, we show the performance quality of our CCP method on Chinese data on the Yi 6b model. We see that, as for other English models, our method outperforms the alternatives over the entire generation length in terms of ROC-AUC.
Appendix F Resources and Expenses
For a single run of data generation and UQ methods evaluation, we spent 12 days of Nvidia A100 GPU compute time. OpenAI API was used mainly for splitting and matching atomic claims, the total cost for all English models (Mistral 7b, Vicuna 13b, Jais 13b, GPT-3.5-turbo) was \$40. The cost of generating English biographies with GPT-3.5-turbo was \$13.
@@ Appendix G Examples of Results Obtained with CCP and the Maximal Probability Baseline
Figures 10, 11 and 12 show examples of the Claim-Conditioned Probability method compared with Maximum Probability. On all examples, our CCP method correctly detects the wrong facts, whereas Maximum Probability mistakenly highlights correct facts.
@@
Figure 10: The Maximum Probability baseline vs Claim-Conditioned Probability (ours).
@@
Figure 11: The Maximum Probability baseline vs Claim-Conditioned Probability (ours).
@@
Figure 12: The Maximum Probability baseline vs Claim-Conditioned Probability (ours).